OpenStudy (anonymous):

if z-7=14. thenz+2 is equal to what?

OpenStudy (anonymous):

z = 14 + 7 z = 21 so 21 + 2 = 23

OpenStudy (dumbcow):

z=21 so 21+2 = 23